A reminder factor theorem find the value of k if on dividing 2 x cube + 3 x square minus kx + 5 by x minus 2 leaves a remainder 7

1
Expert's answer
2020-08-17T18:59:22-0400

When f(x) is divided by x-2, the remainder is 7.


Therefore, using remainder factor theorem, f(2)=7


f(2) = 2(2)3 + 3(2)2- k(2) + 5


7 = 16 + 12 - 2k + 5


7 = 33 - 2k


-2k= -26


k = 13